# Early warning signals of the tipping point in strongly interacting Rydberg atoms

### v1

@@
22footnotetext: J.Z, Z.K.L, L.H.Z contribute equally to this work.11footnotetext: dds@ustc.edu.cn
@@
Early warning signals of the tipping point in strongly interacting Rydberg atoms
Jun Zhang1,2,†    Zong-Kai Liu1,2,†    Li-Hua Zhang1,2,†    Bang Liu1,2    Zheng-Yuan Zhang1,2    Shi-Yao Shao1,2    Qing Li1,2    Han-Chao Chen1,2    Yu Ma1,2    Tian-Yu Han1,2    Qi-Feng Wang1,2    C. Stuart Adams3    Dong-Sheng Ding1,2,⋆    Bao-Sen Shi1,2 1Key Laboratory of Quantum Information, University of Science and Technology of China, Hefei, Anhui 230026, China. 2Synergetic Innovation Center of Quantum Information and Quantum Physics, University of Science and Technology of China, Hefei, Anhui 230026, China. 3Department of Physics, Joint Quantum Centre (JQC) Durham-Newcastle,
@@
(May 1, 2024)
The identification of tipping points is essential for prediction of collapses or other sudden changes in complex systems. Applications include studies of ecology, thermodynamics, climatology, and epidemiology. However, detecting early signs of proximity to a tipping is made challenging by complexity and non-linearity. Strongly interacting Rydberg atom gases offer a model systems that offer both complexity and non-linearity, including phase transition and critical slowing down. Here, via an external probe we observe prior warning of the proximity of a phase transition of Rydberg thermal gases. This warning signal is manifested as a cessation of the variance growth with increasing probe intensity. We also observed the dynamics of the critical slowing down behavior versus different time scales, driving intensities, and atomic densities, thus providing insights into the study of a Rydberg atom system’s critical behavior. Our experiment suggests that the full critical slowing down dynamics of strongly-interacting Rydberg atoms can be probed systematically, thus providing a benchmark with which to identify critical phenomena in quantum many-body systems.
As a result of their inherent interconnectedness and nonlinear dynamics, complex systems have tipping points at which small changes in a specific component or parameter can trigger a disproportionately large response in the entire system [1, 2]. Understanding and identifying these tipping points are essential steps in enabling complex systems to be managed effectively [3, 4, 5, 6, 7]. Effective management could allow us to anticipate and respond to potential tipping points and to implement strategies that can avoid or mitigate their detrimental effects [3, 8, 9]. Furthermore, complex systems often show self-organization [10, 11] and emergent properties [12], which means that their behavior at the collective level cannot be predicted solely based on an understanding of the individual components. As the interactions and dependencies between the components become stronger near the tipping point, the system will then become more susceptible to abrupt shifts [13, 14, 15]. Various factors can contribute to the existence of tipping points in different complex systems; examples include epileptic seizures in medicine [16, 17], systemic market crashes in financial networks [18, 19], climate in nature [20, 21, 22], and phase transitions in physics [23, 24, 25]. Several significant works in the literature have reported on building of theoretical models to predict tipping points [4, 26, 27], with examples that include landscape-flux theory [28, 29], deep learning methods [30, 31], and the critical slowing down method [21, 32, 22], in which early warning signals are introduced.
The long-range interactions that occur between Rydberg atoms provide us with the ability to simulate large scale systems and study non-equilibrium phase transitions, self-organized criticality, and other phenomena [33, 34, 10, 35, 36, 37]. Rydberg atoms capture the interconnectedness between the different components in a system and also provide strong nonlinearity, which is particularly relevant for complex systems [38, 39, 40, 41, 42]. Therefore, the study of Rydberg atom systems aids in understanding of emergent behaviors and also provides insights into the dynamics of complex systems [43, 44, 45, 46, 47, 48]. However, the complex interconnectedness and nonlinearity, the nonstationary dynamics, and the uncertainty evolution of such systems mean that they are difficult to control and manipulate [49, 50, 51, 52]. For example, even small changes in the system parameters can lead to significant responses within the entire atomic system, where the behavior of one component can have far-reaching effects on the overall system dynamics, including inducing rich emergent properties and self-organization dynamics [10]. This also represents a widespread problem among other complex systems. However, no experiments on prediction of the tipping points in Rydberg atom systems have been reported to date, and these systems are worthy of study.
Figure 1: Physical model and early warning signals of the tipping point. (a) A phase transition in a complex network, in which phase 1 and phase 2 represent the different states of matter below and above the tipping point, respectively. These two phases show different properties, and we use the interconnectedness link here between the particles in phase 2, as compared with no available link in phase 1. (b) Two-level atom model including the ground state $\left|g\right\rangle$ and the Rydberg state $\left|r_{1}\right\rangle$, along with the spontaneous radiation rate $\Gamma_{1}$. The atoms are driven from the ground state to the Rydberg state by a laser with a Rabi frequency $\Omega_{1}$ and detuning $\Delta_{1}$. The atoms are then driven by a modulated microwave (MW) electric field with a Rabi frequency of $\Omega_{2}$. The Rydberg energy level is then modified by the many-body interaction strength $V={C_{6}}/{R^{6}}$ [where $C_{6}$ is the van der Waals coefficient and $R$ is the distance between the Rydberg atoms]. (c) Phase transition of the Rydberg atom system. The Rydberg atom population $\rho_{r_{1}r_{1}}$ shows a sudden change versus the square of the Rabi frequency ${\Omega_{1}}^{2}$, where ${\Omega}^{2}_{1,c}$ corresponds to the tipping point. (d) Measured transmission of the probe beam versus the parameter ${\Omega_{p}}^{2}$ under the MW field driving condition. (e) Measured variance from (d). In this process, there is an early warning signal regime, which is marked as the blue shaded area. The variance in this regime is different to the corresponding variance far away from the tipping point, manifesting as a nonlinear growth property. The black dotted line is a linear curve, which is included as a guide for the eye. We fitted the trend of the variance using a solid red line, where the fit function is $y=p(x-20)e^{q(x-20)^{\alpha}}$+$s$ ($p$ = -4.44$\times$10-7 $\pm$ -4.16$\times$10-10, $q$ = -3.89$\times$10-9 $\pm$ 3.72$\times$10-10, $s$ = 2.01$\times$10-6 $\pm$ 1.41$\times$10-9, $\alpha$ = 8.50 $\pm$ 0.04).
Here, we report an experiment in which the tipping point of interacting Rydberg atoms can be predicted by driving the system using a resonant microwave (MW) electric field. We observe that early warning signals are manifested as a symptom regime in which the variance in the transmission no longer grows linearly. By using a nondestructive measurement method with the electromagnetically induced transparency (EIT) detection method, we can coherently map the status of the Rydberg atoms with respect to the probe field. The two metastable states of the non-equilibrium system will display different responses because the non-interacting Rydberg atoms demonstrate relatively small dissipation when compared with that of the interacting Rydberg atoms. Interactions make the Rydberg atoms difficult to synchronize under the driving MW electric field. Therefore, in the experiments, we can identify the point at which the system undergoes a phase transition directly. This work provides predictive capabilities with regard to the behavior of complex systems and will enable better early responses to potential tipping points.
Results
Warning signal model
The inherent interconnectedness and the nonlinear interactions within complex systems make it difficult to predict sudden changes in these systems. Modeling of complex systems can help us to find early warning signals before these sudden changes, as illustrated in Fig. 1(a). The Rydberg atom system represents a good platform for study of complex systems because of the strong interactions that occur between the Rydberg atoms, and there is a non-equilibrium phase transition for Rydberg atoms from the non-interacting phase (NI phase) to the interacting phase (I phase), as illustrated in Fig. 1(b,c).
The Rydberg atom system consists of $N$ interacting three-level atoms with a ground state $\left|g\right\rangle$, a Rydberg state $\left|r_{1}\right\rangle$ with a decay rate of $\Gamma_{1}$, and a Rydberg state $\left|r_{2}\right\rangle$ with a decay rate of $\Gamma_{2}$, as depicted in Fig. 1(b); further details about the experiment can be found in the Methods section. A two-photon process excites the atom from the ground state $\left|g\right\rangle$ to the Rydberg state $\left|r_{1}\right\rangle$ with a Rabi frequency of $\Omega_{1}$ and detuning of $\Delta_{1}$. The intermediate energy level of this two-photon process is eliminated adiabatically and is not shown in Fig. 1(b).
A dual-tone MW field drives the atoms between the Rydberg states $\left|r_{1}\right\rangle$ and $\left|r_{2}\right\rangle$, and this MW field takes the form of $E\rm_{MW}$$\cos(\omega_{1}t)+E\rm_{MW}\cos{[(\omega_{1}+\delta f)t]}$, where $E\rm_{MW}$ represents the amplitude, $\omega_{1}$ is the resonant frequency with respect to the transition between the two Rydberg states $\left|r_{1}\right\rangle$ and $\left|r_{2}\right\rangle$, and $\delta f$ is the MW frequency detuning from resonance. After application of the rotation wave approximation, this dual-tone MW field turns into a periodic driving field with frequency $\delta f$. When the MW field is off and the intensity profile of the probe laser is scanned, a non-equilibrium phase transition occurs that is characterized by a sudden increase in the population of the Rydberg atoms $\rho_{r1r1}$, as shown in Fig. 1(c). When the dual-tone MW field is on, an oscillation effect occurs in the transmission spectrum. The MW field serves as a perturbation in this case and the spectrum represents the response of the Rydberg atoms to this perturbation, as illustrated in Fig. 1(d). When the Rabi frequency of the probe field is low and the system is outside the warning regime, which is shaded in blue in Fig. 1(d), the Rydberg atom system is then a linear system and its response to the MW field is also linear, as indicated by the linear growth observed in the variance as the square of the Rabi frequency $\Omega_{p}^{2}$ increases; see the Methods section for the definition of the variance. However, near the sharp jump in the transmission, there is a warning regime region colored blue, where the variance in the transmission decreases as $\Omega_{p}^{2}$ increases, as shown in Fig. 1(e). This variance, which deviates from the previously linear increase, represents a warning signal.
@@ Results
where $\hat{H}$ is the Hamiltonian and $\mathcal{L}_{\{r_{1},r_{2}\}}$ is the Lindblad jump operator. The Hamiltonian of the system after adiabatic elimination of the intermediate energy level between $\left|g\right\rangle$ and $\left|r\right\rangle$ is given by
@@
where $\sigma_{i}^{gr}$ ($r={r_{1},r_{2}}$) represents the $i$-th atom transition between the ground state $\left|g\right\rangle$ and the Rydberg state $\left|r\right\rangle$; $n_{i}^{r_{1},r_{2}}$ are the population operators of the two Rydberg energy levels $\left|r_{1}\right\rangle$ and $\left|r_{2}\right\rangle$; $V_{ij}^{r_{1}r_{2}}$ and $V_{ij}^{r_{2}r_{1}}$ represent the interactions between the Rydberg atoms in states $\left|r_{1,2}\right\rangle$ and $\left|r_{2,1}\right\rangle$, respectively; and $V_{ij}^{r_{1}r_{1}}$ and $V_{ij}^{r_{2}r_{2}}$ represent the interactions between the Rydberg atoms in states $\left|r_{1}\right\rangle$ and $\left|r_{2}\right\rangle$, respectively.
@@
$\mathcal{L}_{r}=(\gamma_{r}/2)\sum_{i}(\hat{\sigma}_{i}^{rg}\hat{\rho}\hat{\sigma}^{gr}_{i}-\{\hat{n}_{i}^{r},\hat{\rho}\}),$ (2)
@@
where these terms represent the decay process from the Rydberg state $\left|r\right\rangle$ to the ground state $\left|g\right\rangle$.

### v2

@@
Abstract
The identification of tipping points is essential for prediction of collapses or other sudden changes in complex systems. Applications include studies of ecology, thermodynamics, climatology, and epidemiology. However, detecting early signs of proximity to a tipping is made challenging by complexity and non-linearity. Strongly interacting Rydberg atom gases offer model systems that offer both complexity and non-linearity, including phase transition and critical slowing down. Here, via an external probe we observe prior warning of the proximity of a phase transition of Rydberg thermal gases. This warning signal is manifested as a deviation from linear growth of the variance with increasing probe intensity. We also observed the dynamics of the critical slowing down behavior versus different time scales, and atomic densities, thus providing insights into the study of a Rydberg atom system’s critical behavior. Our experiment suggests that the full critical slowing down dynamics of strongly-interacting Rydberg atoms can be probed systematically, thus providing a benchmark with which to identify critical phenomena in quantum many-body systems.
As a result of their inherent interconnectedness and nonlinear dynamics, complex systems have tipping points at which small changes in a specific component or parameter can trigger a disproportionately large response in the entire system [1, 2]. Discovering early warning signals shown in Fig. 1(a) before these tipping points are essential steps in enabling complex systems to be managed effectively, which allows us to implement strategies that can avoid or mitigate the detrimental effect of abrupt changes[3, 4, 5, 6, 7, 8, 9]. Furthermore, complex systems often show self-organization [10, 11] and emergent properties [12], which means that their behavior at the collective level cannot be predicted solely based on an understanding of the individual components. As the interactions and dependencies between the components become stronger near the tipping point, the system will then become more susceptible to abrupt shifts [13, 14, 15]. Various factors can contribute to the existence of tipping points in different complex systems; examples include epileptic seizures in medicine [16, 17], systemic market crashes in financial networks [18, 19], climate in nature [20, 21, 22], and phase transitions in physics [23, 24, 25]. Several significant works in the literature have reported on building of theoretical models to predict tipping points [4, 26, 27], with examples that include landscape-flux theory [28, 29], deep learning methods [30, 31], and the critical slowing down [21, 32, 22] which describes how the dynamics of a system change as it approaches a tipping point, by these ways early warning signals are introduced.
The long-range interactions that occur between Rydberg atoms provide us with the ability to simulate large scale systems and study non-equilibrium phase transitions, self-organized criticality, and other phenomena [33, 34, 10, 35, 36, 37]. The connection between Rydberg atom systems and complex systems lies in their shared features of the complex interconnectedness, strong nonlinearity and emergent behaviors [38, 39, 40, 41, 42, 43, 44, 45]. Therefore, the study of Rydberg atom systems aids in understanding of emergent behaviors and also provides insights into the dynamics of complex systems [46, 47, 48, 49, 50, 51]. However, the complex interconnectedness and nonlinearity, the nonstationary dynamics, and the uncertainty evolution of such systems mean that they are difficult to control and manipulate [52, 53, 54, 55]. For example, even small changes in the system parameters can lead to significant responses within the entire atomic system, where the behavior of one component can have far-reaching effects on the overall system dynamics, including inducing rich emergent properties and self-organization dynamics [10]. However, no experiments on prediction of the tipping points in Rydberg atom systems have been reported to date, and these systems are worthy of study.
Figure 1: Physical model and early warning signals of the tipping point. (a) A phase transition in a complex system, in which phase 1 and phase 2 represent the different states of matter below and above the tipping point. (b) Two-level atom model including the ground state $\left|g\right\rangle$ and the Rydberg state $\left|r_{1}\right\rangle$, and the atoms are driven by a dual-tone microwave (MW) field. (c) Phase transition of the Rydberg atom system. (d) Measured transmission of the probe beam versus the parameter $\Omega_{p}^{2}$ under the MW field driving condition. (e) Measured variance from (d). The early warning signal regime manifest a nonlinear growth property, which is marked as the blue shaded area. The black dotted line is a linear curve. We fitted the trend of the variance using a solid red line, where the fit function is $y=p(x-20)e^{q(x-20)^{\alpha}}$+$s$ ($p$ = -4.44$\times$10-7 $\pm$ -4.16$\times$10-10, $q$ = -3.89$\times$10-9 $\pm$ 3.72$\times$10-10, $s$ = 2.01$\times$10-6 $\pm$ 1.41$\times$10-9, $\alpha$ = 8.50 $\pm$ 0.04).
Here, we report an experiment in which the tipping point of interacting Rydberg atoms can be predicted by driving the system using a resonant microwave (MW) electric field. We observe that early warning signals are manifested as a symptom regime in which the variance in the transmission no longer grows linearly. In our system, the signals are extracted from the system’s response to MW, whereas the signals in other complex systems are extracted from internal fluctuations[30, 31, 16]. By using a nondestructive measurement method with the electromagnetically induced transparency (EIT), we can coherently map the status of the Rydberg atoms with respect to the probe field. The two metastable states of the non-equilibrium system display different responses because the low-density Rydberg atoms (phase 1) demonstrate relatively small dissipation when compared with that of the high-density Rydberg atoms (phase 2). Interactions make the Rydberg atoms difficult to synchronize under the driving MW field. Therefore, in experiment, we can identify the tipping point at which the system undergoes a phase transition directly. This work provides predictive capabilities with regard to the behavior of complex systems and will enable better early responses to potential tipping points.
Warning signal model. The inherent interconnectedness and the nonlinear interactions within complex systems make it difficult to predict sudden changes in these systems. Modeling of complex systems can help us to find early warning signals before these sudden changes, as illustrated in Fig. 1(a). The Rydberg atom system represents a good platform for study of complex systems because of the strong interactions that occur between the Rydberg atoms, and there is a non-equilibrium phase transition for Rydberg atoms from the phase 1 to the phase 2, as illustrated in Fig. 1(b, c).
The Rydberg atom system consists of $N$ interacting three-level 85-Rubidium (85Rb) atoms with a ground state $\left|g\right\rangle$, a Rydberg state $\left|r_{1}\right\rangle=\left|50D_{5/2}\right\rangle$ with a decay rate of $\Gamma_{1}$, and a Rydberg state $\left|r_{2}\right\rangle=\left|51P_{3/2}\right\rangle$ with a decay rate of $\Gamma_{2}$, as depicted in Fig. 1(b). The atoms are driven into the Rydberg state $\left|r_{1}\right\rangle$ using a probe beam and a coupling beam. The probe beam excites atoms from the ground state $|g\rangle=|5S_{1/2},F=3\rangle$ to the intermediate excited state $|e\rangle=|5P_{3/2},F=4\rangle$, and the coupling beam drives the transition from $|e\rangle$ to the Rydberg state $|r_{1}\rangle$. As depicted in Fig. 1(b), the intermediate energy level $|e\rangle$ is adiabatically eliminated, thus effective coupling is between the ground state $|g\rangle$ and the Rydberg state $|r_{1}\rangle$ with a Rabi frequency $\Omega_{1}$ and a detuning $\Delta_{1}$. The beams counterpropagate through a thermal Rb vapor, and the EIT spectrum is obtained by detecting the probe beam intensity. Further details about the experiment can be found in the Supplemental Material.
A dual-tone MW field drives the atoms between the Rydberg states $\left|r_{1}\right\rangle$ and $\left|r_{2}\right\rangle$, and this MW field takes the form of $E\rm_{MW}$$\cos(\omega_{1}t)+E\rm_{MW}\cos{[(\omega_{1}+\delta\omega)t]}$, where $E\rm_{MW}=\hbar\Omega_{2}/\mu\rm_{MW}$ represents the MW field amplitude [$\Omega_{2}$ is Rabi frequency of the MW field and $\mu\rm_{MW}$ is the transition dipole moment], $\omega_{1}$ is the resonant frequency with respect to the transition between the two Rydberg states $\left|r_{1}\right\rangle$ and $\left|r_{2}\right\rangle$, and $\delta\omega$ is the MW frequency detuning from resonance. After application of the rotation wave approximation, this dual-tone MW field turns into a periodic driving field with frequency $\delta\omega$.
When the MW field is off and the intensity profile of the probe laser is scanned, a non-equilibrium phase transition occurs that is characterized by a sudden increase in the population of the Rydberg atoms $\rho_{r1r1}$, as shown in Fig. 1(c). When the dual-tone MW field is on, an oscillation effect occurs in the transmission spectrum. The MW field serves as a driving in this case and the spectrum represents the response of the Rydberg atoms to this driving, as illustrated in Fig. 1(d) .
To acquire the warning signals before the tipping point, we calculate the variance using the moving window algorithm. The formula is
$Var=\frac{1}{m-1}\sum_{i=1}^{m}\left|a_{i}-\mu\right|^{2}$ (1)
where $a_{i}$ is the transmission intensity of probe beam, $m$ is the data length of a window, and $\mu$ is the mean value in the window.
When the Rabi frequency of the probe field $\Omega_{p}$ is low and the system is outside the warning regime, which is shaded in blue in Fig. 1(d), the Rydberg atom system is a linear system and its response to the MW field is also linear, as indicated by the linear growth observed in the variance as the square of the Rabi frequency $\Omega_{p}^{2}$ increases. However, near the sharp jump in the transmission, there is a warning regime region colored blue. The variance in the region deviates from the previously linear increase, represents a warning signal.
To model the above process, we consider the effect of interactions-induced dissipation, in which interactions between distinct Rydberg atoms could cause spatially dependent phase shifts and increases the total relaxation rate [39, 10]. The detuning and the decay terms are replaced by $\Delta_{1}\rightarrow\Delta_{1}-V\rho_{r_{1}r_{1}}^{\beta}$ and $\Gamma_{r_{1}}\rightarrow\Gamma_{r_{1}}-V\rho_{r_{1}r_{1}}^{\epsilon}$ [23, 10, 34, 39, 56] in the mean-field approximation, where $\beta$ and $\epsilon$ are exponent factors. By solving the master equation and expanding the solution $\rho_{r_{1}r_{1}}$ as a function of $\Omega_{1}$ to the second order, the transmission against $(\Omega_{1}/2\pi)^{2}$ is obtained. Here, $\beta=1$ and $\epsilon=3$ are selected to model the results better.
To observe the full dynamics near the tipping point, we fit the trend of variance using the function $y=p(x-20)e^{q(x-20)^{\alpha}}$+$s$ [see the solid red line in Fig. 1(e)], where $p$, $q$ and $s$ are the coefficients, and $\alpha$ represents the critical exponent which describes the features of the critical slowing down in the non-equilibrium system. The critical exponent $\alpha$ characterizes the system’s behavior near the tipping point and describes how it scales as system approaches the tipping point.
Figure 2: Theoretical and experimental results without and with the driving MW field. (a), (b) Theoretical simulations of the population $\rho_{r_{1}r_{1}}$ of the Rydberg atoms with different electric field amplitudes: (a) $E_{\mathrm{MW}}$ = 0 $\mathrm{mV/cm}$ and (b) $E_{\mathrm{MW}}$neq 0 $\mathrm{mV/cm}$. (c) Variance calculated from (b). (d), (e) Experimental transmission spectra under the different MW field amplitudes: (d) $E\rm_{MW}$ = 0 $\mathrm{mV/cm}$ and (e) $E\rm_{MW}$ = 1.64 $\mathrm{mV/cm}$. (f) Measured variance from (e).
Transmission spectra under the dual-tone MW field driving conditions. We apply the dual-tone MW field to the theoretical model and simulate the transmission, where the MW field takes the form of $\Omega_{2}=\mu\rm_{MW}E_{\mathrm{MW}}(1+\mathrm{Cos}(\delta{\omega}t))/\hbar$. The simulated results for $\rho_{r_{1}r_{1}}$ versus $(\Omega_{1}/2\pi)^{2}$ are shown in Fig. 2(a-b), where panel (a) corresponds to the case of the phase transition without the MW field ($E_{\mathrm{MW}}=0$) and panel (b) shows the case with the dual-tone MW field ($E_{\mathrm{MW}}\neq 0$). Fig. 2(c) shows the variance that was calculated from Fig. 2(b), with nonlinear behavior in the region where 2.2 ($2.2\text{ }(\mathrm{MHz})^{2}<(\Omega_{1}/2\pi)^{2}<3.5\text{ }(\mathrm{MHz})^{2}$. This nonlinear behavior serves as the early warning signals, but the signals are difficult to extract without the MW field, shown in Figs. 2(a) and (d). Applying the dual-tone MW field amplifies the signals and makes the interactions-induced dissipation easier to detect.
By scanning the probe field intensity, we can observe the non-equilibrium phase transition in the Rydberg atom system, as illustrated in Fig. 2(d). To subtract the early warning signals before the tipping point is reached, we experimentally apply a dual-tone MW electric field to drive the transition between the two different Rydberg states $|50D_{5/2}\rangle$ and $|51P_{3/2}\rangle$, with a center frequency of $2\pi\times 17.041$ GHz and a frequency difference of $\delta\omega$=$2\pi\times 0.2$ MHz. The MW field induces a small ac Stark shift that moves the tipping point and varies the Rydberg atom population. The MW field causes the tipping point to move forward when compared with the case without the MW field, and the oscillation appears in the transmission spectrum in Fig. 2(e). By calculating the variance from the oscillated transmission, we observe a nonlinear variance when $(\Omega_{p}/2\pi)^{2}>26\text{ }(\mathrm{MHz})^{2}$, indicating the existence of an early warning signal, as shown in Fig. 2(f). The experimental results show good agreement with the expectations of the theoretical simulations.
From the results shown in Fig. 2(f), the measured variance reduces nonlinearly when $(\Omega_{p}/2\pi)^{2}>$ 27.5 $(\mathrm{MHz})^{2}$. The increased $\Omega_{p}$ adds a number of excited Rydberg atoms, which makes the system dimension higher and then induces a strong critical slowing down effect. In a relatively low-dimensional system where the number of degrees of freedom is limited, localized interactions between neighboring Rydberg atoms are dominant. As a result, the critical slowing down process in a low-dimensional system is primarily influenced by these localized interactions. In this case, all the excited Rydberg atoms in the system can be synchronized rapidly, thus allowing faster recovery from the perturbation. In contrast, a high-dimensional system has large numbers of interacting Rydberg atoms, which results in a rich interconnectedness within the system. In this system, the critical slowing down effects are amplified because of the extended reach and the influence of the long-range interactions. Perturbations take longer times to propagate through the system, and the process of recovery from disturbances thus becomes slower.
Transmission spectra versus scanning time. The non-equilibrium phase transition in the Rydberg atom system can be modeled using a double-well transition [33], with each well standing for one bi-stable state, as depicted in Fig. 3(a)-(c). The dissipation of phase 2 is greater than that of phase 1 because of the interactions among the Rydberg atoms.
When the population of Rydberg atoms is increased by $\Omega_{p}^{2}$, a phase transition occurs from the left well to the right well that crosses the barrier of the double well, as illustrated in Fig. 3(b). Near the tipping point $\Omega_{p,c}^{2}$, the critical slowing down phenomenon occurs; this term refers to an effect where the system’s response time to the driving MW field becomes slower (here, the MW field is regarded as a perturbation to the system). After the phase transition, the system is then in the phase 2, and the strong interactions between the Rydberg atoms make the system’s dissipation so large that the variance drops to a very low level.
A fast scan accumulates a small number of Rydberg atoms because of the small interval time. This causes the system to have low dissipation until the phase transition occurs, which means that the warning regime is almost invisible [see Fig. 3(d) and Fig. 3(e) for the scanning times of Ts= 0.25 ms and Ts= 0.5 ms, respectively]. As the light intensity increases, the system’s response to the MW field follows the increasing light intensity well.
Figure 4: Variance measured at different temperatures. We measured the different variance characteristics at temperatures of (a) T=47.2C, (b) T=47.7C, (c) T=48.2C, and (d) T=49.1C. The early warning signal regimes shown in (a) and (d) are marked as the blue shaded areas for comparison. The solid red lines are the fittings with the function $y=p(x-20)e^{q(x-20)^{\alpha}}$+$s$, where $\alpha$ = 16.46 $\pm$ 0.22 for T=47.2C, $\alpha$ = 12.69$\pm$ 0.10 for T=47.7C, $\alpha$ = 10.89 $\pm$ 0.07 for T=48.2C, and $\alpha$ = 8.59 $\pm$ 0.07 for T=49.2C.
Transmission spectra at different temperatures. In our experiments, we also studied how the interaction between the Rydberg atoms affects the warning signals. A higher temperature leads to a higher interaction level, and vice versa. We changed the atomic density by varying the temperature and measured the variance of the oscillated transmission. Figure 4(a)-(d) show the measured results. We found that the variance before the phase transition was decreasing gradually as the temperature increased. For comparison, we define the width of the warning regime as $\delta\Omega_{\mathrm{p}}^{2}$. We found that the early warning signal regime shown in Fig. 4(d) was broader than that shown in Fig. 4(a), as given by ($\delta\Omega_{\mathrm{p}}/2\pi)^{2}=1.5\text{ }(\mathrm{MHz})^{2}$ at T = 47.2C and ($\delta\Omega_{\mathrm{p}}/2\pi)^{2}=3.0\text{ }(\mathrm{MHz})^{2}$ at T = 49.1C. Therefore, it is easier to identify the early warning signal at higher temperatures.This is explained by the fact that the interaction-induced dissipation causes the system to be difficult to synchronize when driven by the MW field, and it then leaves the linear response regime. This enhanced critical slowing down process in high-dimensional systems can be attributed to the increased system complexity and the interdependences among the Rydberg atoms.
The early warning signals of the Rydberg atom systems under different spatial structure conditions and scales [from their distinct densities] can also be probed, and the measured exponent $\alpha$ varies with the temperature [e.g., for T = 47.2C, $\alpha$ = 16.46 $\pm$ 0.22 in Fig. 4(a), but $\alpha$ = 8.59 $\pm$ 0.07 for T = 49.2C in Fig. 4(d)], thus indicating distinct critical dynamics before occurrence of a phase transition. This confirms that the early warning signal arises from the interactions between the Rydberg atoms, even when the system is still in the phase 1. This finding is consistent with the fact that the non-equilibrium phase transition in the Rydberg atom system occurs because of the interactions among a few local atoms, which eventually spreads to involve the atoms on a global scale. Using the MW field, we can detect small numbers of locally interacting Rydberg atoms, even without causing a phase transition to occur in the system.
Discussion. It should be noted that the measured variance in our experiment was reduced in the early warning regime, which differs from the increased variance that was reported previously [30, 31, 16]. In those cases, the variance reveals the fluctuations of internal variables of the system; while in our system, the variance represents the strength of oscillation under external MW field driving. In this way, the interaction-induced dissipation before the tipping point is related to the amplitude of oscillation, thus allowing us to characterize the full dynamics of the critical slowing down process. The early warning signals for the phase transition cause us to conclude that the interaction-induced dissipation appears first, and the phase transition occurs after that. This is also consistent with the results from our mean-field method, in which the exponent factor $\epsilon$ was greater than $\beta$.
We acknowledge funding from the National Key R&D Program of China (Grant No. 2022YFA1404002), the National Natural Science Foundation of China (Grant Nos. U20A20218, 61525504, and 61435011), the Anhui Initiative in Quantum Information Technologies (Grant No. AHY020200), the major science and technology projects in Anhui Province (Grant No. 202203a13010001).
J.Z implemented the physical experiments with L.H.Z. D-S.D. and J.Z conducted the theoretical model. All authors contributed to the experiment preparation and data discussions regarding the results. D-S.D. conceived the idea and support this project.
@@ Abstract
Supplemental Material for “Early warning signals of the tipping point in strongly interacting Rydberg atoms”
Jun Zhang1,2,†, Li-Hua Zhang1,2,†, Bang Liu1,2, Zheng-Yuan Zhang1,2, Shi-Yao Shao1,2,
Qing Li1,2, Han-Chao Chen1,2, Zong-Kai Liu1,2, Yu Ma1,2, Tian-Yu Han1,2,
Qi-Feng Wang1,2, C. Stuart Adams3, Bao-Sen Shi1,2, Dong-Sheng Ding1,2,⋆
1Key Laboratory of Quantum Information, University of Science and Technology of China,
Hefei, Anhui 230026, China.
2Synergetic Innovation Center of Quantum Information and Quantum Physics,
University of Science and Technology of China, Hefei, Anhui 230026, China.
3Department of Physics, Joint Quantum Centre (JQC) Durham-Newcastle,
@@
22footnotetext: J.Z, L.H.Z contribute equally to this work.11footnotetext: dds@ustc.edu.cn
Theoretical model
@@
where $\hat{H}$ is the Hamiltonian and $\mathcal{L}_{\{r_{1},r_{2}\}}$ is the Lindblad jump operator. The Hamiltonian of the system after adiabatic elimination of the intermediate energy level between $\left|g\right\rangle$ and $\left|r_{1}\right\rangle$ is given by
@@
where $\sigma_{i}^{gr_{1}}$ and $\sigma_{i}^{r_{1}r_{2}}$ represent the $i$-th atom transition between $\left|g\right\rangle$ and $\left|r_{1}\right\rangle$ and $\left|r_{1}\right\rangle$ and $\left|r_{2}\right\rangle$; $n_{i}^{r_{1},r_{2}}$ are the population operators of the two Rydberg energy levels $\left|r_{1}\right\rangle$ and $\left|r_{2}\right\rangle$; $V_{ij}^{r_{1}r_{2}}$ and $V_{ij}^{r_{2}r_{1}}$ represent the interactions between the Rydberg atoms in states $\left|r_{1,2}\right\rangle$ and $\left|r_{2,1}\right\rangle$, respectively; and $V_{ij}^{r_{1}r_{1}}$ and $V_{ij}^{r_{2}r_{2}}$ represent the interactions between the Rydberg atoms in states $\left|r_{1}\right\rangle$ and $\left|r_{2}\right\rangle$, respectively.
@@
$\mathcal{L}_{r_{1}}=(\Gamma_{1}/2)\sum_{i}(2\hat{\sigma}_{i}^{r_{1}g}\hat{\rho}\hat{\sigma}^{gr_{1}}_{i}-\{\hat{n}_{i}^{r_{1}},\hat{\rho}\}),$ (3)
$\mathcal{L}_{r_{2}}=(\Gamma_{2}/2)\sum_{i}(2\hat{\sigma}_{i}^{r_{2}r_{1}}\hat{\rho}\hat{\sigma}^{r_{1}r_{2}}_{i}-\{\hat{n}_{i}^{r_{2}},\hat{\rho}\}),$ (4)
@@
where these terms represent the decay process from the Rydberg state $\left|r_{1}\right\rangle$ to the ground state $\left|g\right\rangle$, and the Rydberg state $\left|r_{2}\right\rangle$ to the Rydberg state $\left|r_{1}\right\rangle$.